### v1

# Curiosity Driven Exploration to Optimize Structure-Property Learning in Microscopy

Abstract
I Abstract
Rapidly determining structure-property correlations in materials is an important challenge in better understanding fundamental mechanisms and greatly assists in materials design. In microscopy, imaging data provides a direct measurement of the local structure, while spectroscopic measurements provide relevant functional property information. Deep kernel active learning approaches have been utilized to rapidly map local structure to functional properties in microscopy experiments, but are computationally expensive for multi-dimensional and correlated output spaces. Here, we present an alternative lightweight curiosity algorithm which actively samples regions with unexplored structure-property relations, utilizing a deep-learning based surrogate model for error prediction. We show that the algorithm outperforms random sampling for predicting properties from structures, and provides a convenient tool for efficient mapping of structure-property relationships in materials science.
II Introduction
Determining structure-property relationships is crucial to the development of new materials with desired functional properties, and therefore rapid determination is critical to accelerate material design and optimization. More generally, in the context of autonomous and self-driving laboratories, rapidly determining the relevant relationships between structure and function is critical to optimizing relevant chemical synthetic and processing pathways for molecular and materials optimization and discoverytomsdl2024 .
Microscopy, in particular scanning probe and electron microscopy, provides a powerful method to locally image structures with nanoscale or atomic resolutionbian2021scanning . In addition, the ability to spatially probe spectroscopic properties allows for correlating the local structure with site-specific functional properties. Traditionally, spatially resolved measurements are performed across a grid of points using techniques such as atomic force microscopy force mapping, scanning tunneling spectroscopy, or electron energy loss spectroscopy in a scanning transmission electron microscope. The downside of this method is that (a) only a small number of points can be probed given a limited experimental time budget, and (b) increasing the number of measured spectroscopic points to increase resolution can result in irreversible tip and/or sample damage. Machine learning applications in scientific methodshase2019next , especially in the past decade, have impacted imaging techniquesmoen2019deep ; archit2025segment ; krull2020artificial ; kalinin2021automated . Adaptive sampling methods based on route optimizationkandel2023demonstration ; godaliyadda2016supervised ; noack2019kriging and sparse samplingcheca2023high ; gura2021spiral have been used for efficient image reconstruction. In particular, with regard to learning structure-property relationships, deep kernel active learning (DKL) approaches have been utilized to adaptively sample material properties using input image patches acquired in the imaging mode on the microscope liu2022experimental . This was shown to be highly efficient in correlating local ferroelectric domain structures with specific features of ferroelectric hysteresis loops in the pioneering work by Liu et al liu2022experimental . That work was subsequently extended to other modalities, including conductive atomic force microscopy, electron microscopy and scanning tunneling microscopy liu2023exploring ; roccapriore2022physics ; narasimha2024multiscale . However, DKL, and indeed all Bayesian optimization approaches, utilize a scalarizer function to reduce high-dimensional spectroscopic measurements to a single scalar quantity that is used as the target for optimizationhickman2025atlas . While this approach is a suitable method to optimize for a given target property, the exploratory power is limited because of the loss of spectroscopic features that are not accounted for by the scalarizer function. Although multi-objective optimization is possible, attempting to develop Gaussian based methods for large output spaces (e.g., above 10 dims) where the outputs are correlated is at present computationally intractable. In principle, ensembles of DKL models for uncorrelated outputs are also a feasible solution, although in practice, spectral outputs tend to be correlated and this strategy is therefore not viable.
Here, we present alternate methods relying on surrogate models of error prediction, which we term curiosity-driven exploration, analogous to the usage of the term in reinforcement learning zheng2021episodic ; thiede2022curiosity ; pathak2017curiosity . These methods are based on standard deep neural networks with an encoder-decoder structure that have been employed in the past to predict spectra from images (Im2spec) and images from spectra (Spec2im)kalinin2021toward . When the goal is to minimize the loss of an Im2spec or Spec2im model, the optimal scalarizer function is difficult, if not impossible to find. As a solution, we instead determine which spectra to measure by training an auxiliary network to predict Im2spec reconstruction error. The curiosity-driven approach involves sampling regions with high values of the predicted error, so as to attempt to rapidly reduce the error of these models.
The paper presents two workflows: The first consists of an ensemble of Im2spec models that is used for spectral prediction, followed by an error model that trains on the spectral mismatch error. In the second method, the error model utilizes the latent space embeddings of an autoencoder to correlate with spectral mismatch. These algorithms, inspired by curiosity-driven reinforcement learning, actively sample spectra for which the structure-property relations have not yet been learned. We first demonstrate and optimize the efficacy of our methods on a pre-acquired dataset. Finally, we implement an algorithm on an atomic force microscope (AFM) to actively learn structure-property relationships in a ferroelectric thin film and discuss possible extensions.
Figure 1: Description of the spectral search method on pre-acquired dataset. (a) Spatial dependence of the spectral property that is correlated with the sample region shown on the left. The patches on the sample image serve as the structural inputs that is correlated to the spectral output. (b) An ensemble of different Im2spec models are trained on an initial set of data. (c) The best model is then used to predict the spectral output corresponding to the training image inputs. (d) Error model consists of the encoder and the latent embedding part of the Im2spec model. This is conjoined with a decoder that is used to train with the spectral mismatch L1-error. The parameters of the encoder (and the latent embedding) are frozen while training the error model. (e) The error model is used to prediction for the image patches across the sample region. The next points are decided using the acquisition function, and incorporated into the training set for the subsequent iteration.
III Results and Discussions
III.1 Im2spec encoded error model.
The structure-property relations described in this work correlate to the ferroelectric response of PbTiO3 samples measured using band excitation piezoresponse microscopy (BE-PFM). We use the structure information from the piezoresponse imaging data acquired from the AFM scanned images, while the property is measured using the spectra collected using band excitation piezoresponse spectroscopy (BEPS) data. This dataset is very similar to one captured and published earlier, and details about the measurement can be found elsewhereslautin2024bayesian .
Fig 1 illustrates the active learning workflow described in this section. Fig 1(a) shows sample dataset which shows spatial dependence of the local structure and its influence on the observed spectrum. Here, the local structure, indicated by the square patch, influences the spectrum measured in that region. We initially start by considering a training set where the inputs are the image patches (each patch of size (16 x 16) pixels) while the outputs are the spectra (256 points) corresponding to each patch. This initial set is used to train Im2spec models.
In this workflow, we use an ensemble of Im2spec models to offer flexibility for variations in the training data, as shown in Fig 1(b). While the models are primarily based on the convolutional networks, variations in the architecture and the hyperparameters have been introduced to enable wider adaptability. A brief description of the encoder architectures used in the model set is provided in Table 1. We designate the size of the latent dimension as three. An initial dataset of the image patches is used to train the Im2spec models. Once trained, the best model of the ensemble is chosen based on the minimum validation loss.
Im2spec model name Encoder architecture
im2spec Convolution block (3 layers, leaky_relu = 0.1, dropout = 0.5)
im2spec_2 Convolution block (3 layers, leaky_relu = 0.2, dropout = 0.1)
im2spec_3 Convolution block (3 layers, leaky_relu = 0.2, dropout = 0.1), Dilated block (4 layers)
im2spec_4 Resnet module (depth = 3), Convolutional block (3 layers, leaky_relu = 0.2, dropout = 0.2)
im2spec_5 Resnet module (depth = 3), Dilated block (4 layers)
Table 1: Encoder architecture of the Im2spec models used in the ensemble
The selected model is then used to predict the spectral output on the image inputs that were previously used for training (as shown in Fig 1(c)). This prediction is compared with the original spectrum, and the mismatch error is assigned to every image within the training set. We use the L1 loss to quantify the spectral mismatch in this method. Fig 1(d) shows the error model that consists of the Im2spec-encoder (which includes the latent embedding layer) conjoined with a different set of decoder layers. During the error model training, the encoder part of the model is frozen while the decoder weights are updated.  The next step involves the error prediction for the entire set of image patches across the sample region, as shown in Fig 1(e). The error predictions are used to compute the acquisition function to determine and sample the next set of spectral points in an iterative active learning fashion.
Our studies show that the best Im2spec model does not change frequently with minor changes in the training data set. Our code allows for the assignment of a probability associated with ensemble training to minimize unnecessary training procedures. In the results described in this section, we perform ensemble training randomly over 20% of the iterations (and the starting iteration). The remaining iterations involve model training using the pre-determined best-Im2spec-model.
The acquisition function used in this method is an empirical equation and is given as:
$A_{j}=1-e^{-\lambda|L_{j}-(1-\beta)|}$
where $L_{j}$ is the L1 error normalized in the range [0, 1]. The $\beta$ parameter controls the rate of exploration and exploitation, while the prefactor $\lambda$ controls the smoothness of the acquisition function (higher $\lambda$ indicates better smoothness). The acquisition function varies monotonically with the error values for $\beta$ = 1 and inversely for $\beta$=0. This allows us to tune sampling from exploitation to exploration as we increase the beta hyperparameter from 0 to 1. In the results described in this section, we study model exploration for different $\beta$ values while maintaining $\lambda$ = 0.1.
The workflow starts with an initial dataset consisting of 245 image-spectrum pairs (20% of the total dataset). Each iteration consists of two model training events - the im2spec ensemble models and the error model. As shown in Fig 1(e), we obtain the prediction of the error values at the end of each iteration. We use the acquisition to sample the next batch of ten points in every iteration. In the results described in this section, we study the model behavior over twenty iterations totaling two hundred explored acquisitions.
Fig 2 shows the workflow results where we test the model for varying values of the $\beta$ parameter in the range [0, 1]. The figures in panel Fig 2(a)-(c) show the the results for $\beta$ = 0, which samples the points associated with low-error values. Fig 2(a) shows the posterior error prediction at the end of iterations. The red scatter points indicate the points of low error values selected by the acquisition function for subsequent sampling. Fig 2(b) shows the points that indicate all acquisitions at the end of the active learning process overlaid on the topographic image. Fig 2(c) shows the statistics of the determined L1 error across all iterations. Given the conservative nature of the exploration associated with $\beta$ = 0, the cumulative error of the points selected by the active learning process is lower than randomly sampled images.
Next, we study the effect of dominant exploration with $\beta$= 1, and the results are shown in Figs. 2(d)-(f). The higher value of the $\beta$ parameter seeks out sample regions with the highest fitting error. This is evident in Fig 2(d), which shows the points selected for sampling in areas with high predicted error. In contrast to the points acquired in Fig 2(b) (for $\beta$ =0), Fig 2(e) shows areas where the points are sampled at complementary locations across the sample region. A closer observation reveals higher sampling in the interfacial regions around the domain wall, where the structure-property correlations are not straightforward. The curious, explorative nature of the model is reflected in Fig 2(f), which shows higher errors associated with the sampled locations.
Figure 2: Results of the error prediction model in active learning on the pre-acquired BEPS data. Panel (a)-(c) shows the results of the model for beta = 0. (a) Error prediction across the sample region. scattered points indicate points for subsequent acquisitions. All scale bars indicate a length of 100 nm. (b) shows the distribution of acquisitions over 20 iterations (200 points) overlayed on topographic image of the sample. (c) shows the distribution of the L1-error for the acquired spectra. Panel (d)-(f) shows the similar results for active learning-based sampling with $\beta$ = 1, while figures (g)-(i) show the results for active learning based sampling for $\beta$ = 0.5.
While we see extreme examples of exploitation and exploration for $\beta$ = 0 and 1, respectively, intermediate values of the $\beta$ can be used to balance exploration and exploitation. Fig (g)-(i) shows the results for $\beta$ = 0.5. This samples regions that correspond to both higher and lower values of the error prediction. This is reflected in the intermediate range of the mean error shown in Fig 2(i).
In an encoder-decoder model, the latent representations determine the efficiency of the reconstruction. We study the latent embeddings that bridge the two different models to gain insights into the workings of the error model and to interpret the essential features that determine the model output. The latent distributions of the model predictions are described in Fig 3 for the active learning process across different values of $\beta$. Fig 3(a)-(c) represents the latent space distributions for $\beta$ = 0. The conservative nature associated with it is seen as the exploration points are localized at the high-density region of the latent space. Further, to correlate the latent encodings to points in the real space, we use a k-means cluster on the latent space, as shown in Fig 3(b), and these are mapped to the real-space coordinates shown in Fig 3(c). A similar analysis is performed for exploration related to $\beta$ = 1, shown in Fig 3(d)-(f). Here, higher exploration has resulted in a dispersed latent distribution. Further exploration points are comparatively sampled in the sparse region of the latent space. In the real space mapping (Fig 3(f)), this translates to acquisition in the complementary areas and corroborates with the data shown in Fig 2(e). However, for intermediate values of the $\beta$ = 0.5, we see sampling across the distribution of the latent space, indicating a balance of exploration and exploitation.
Figure 3: Latent space representations for the active learning-based sampling. Panel (a)-(c) shows the results of the model for beta = 0. (a) Latent distribution of the model prediction for all image-patch inputs. (b) depicts the latent distribution clustered into three components using k-means clustering. (c) Correlation to the real-space coordinates corresponding each of the points in the latent space. Panel (d)-(f) shows the similar results for active learning-based sampling with $\beta$ = 1 while figures (g)-(i) show latent distribution results for active learning-based sampling for $\beta$ = 0.5
These results describe the error prediction methods in conjunction with the acquisition function where the $\beta$ parameter is used to control exploration. The methodology is denoted as a curiosity-driven for $\beta$ = 1 where the model actively searches for the regions of higher predicted error to better learn diverse structure-spectral correlations.
Further, the latent space analysis emphasizes the importance of the latent distributions in its ability to reconstruct the spectra and predict the reconstruction error. These embeddings serve as compressed, structured representations of input data, capturing essential features of the input images. Given this knowledge, in the subsequent section, we implement a generalized methodology to extract latent representations from an autoencoder while efficiently sampling points from the latent space for active learning based acquisitions.
III.2 Autoencoder-based error model
Figure 4: Diagram of Curiosity Algorithm implementation with Im2spec
This section describes the autoencoder based error model. Here we use a similar experimental dataset - PFM-based experimental data on a 200nm (110) $PbTiO_{3}$ thin film sample grown on $SrTiO_{3}$. The structural information contained in the image patches (patch size (11$\times$11 pixels)) with switching spectroscopy spectra (spectrum length:64) captured at low frequency (off-resonance) using an interferometric atomic force microscope from Oxford Instruments (Vero). The baseline curiosity algorithm works as follows: After the sample is imaged an autoencoder is trained on all image patches. Then spectra are acquired on a small number of random initialization points, which are used to train the Im2spec model. The error predictor is then trained on the image patch latent encodings and the Im2spec mean squared error (MSE) for the initial points. Then, a forward pass through the error predictor is performed for all image patch encodings. The spectra of the point with the highest predicted error is then sampled. This continues iteratively where Im2spec and the error predictor are trained on updated training dataset. The overall workflow is illustrated in Fig 4. It should be noted that the error predictor model utilizes dropout to provide an estimate of the uncertainty on the prediction.
This algorithm is sensitive to the initialization points. If the initial data is not representative of the larger distribution, the algorithm is prone to getting stuck in a local minima. The error predictor then poorly estimates the Im2spec error for unrepresented data, and therefore fails to sample certain points optimal for reducing Im2spec loss. Therefore for sparse sampling across the distribution, we train an autoencoder on the image patches and then sample the initialization points that are far apart in the autoencoder’s latent space. One choice is to utilize $k$-means clustering in the latent space, with $k$ equal to the number initialization points. This was followed by choosing the points closest to each respective cluster centroid as the initialization points.
To encourage exploration within the latent representations, we reward points that are far away from previously sampled points in the Im2spec latent space. A natural choice for this exploration reward, $E_{j}$, is the harmonic mean of euclidean distances in the latent space to previously measured points:
$E_{j}=\left(\sum_{\text{measured }i}\frac{1}{|\ell_{i}-\ell_{j}|}\right)^{-1}$
where $\ell_{i}$ denote the Im2spec latent encodings of the image patches. Denoting the error predictions as $C_{j}$, a viable acquisition function, analogous to the epsilon-decreasing strategy for the multi-armed bandit problem slivkins2024introductionmultiarmedbandits , is given by:
$A_{j}=(1-e^{-\lambda n})C_{j}+e^{-\lambda n}E_{j}$
where $n$ is the number of spectra measured so far. Finally, incorporating uncertainty classification in both the error predictor and Im2spec model, and modifying the acquisition function accordingly, would improve exploration. Due to the high dimensional output of Im2spec, we chose to utilize Monte Carlo Dropoutgal2016dropout for uncertainty estimation. For the error predictor, other methods such as deep kernel learning or a fully Bayesian final layer are also feasible. The exploration reward and model uncertainty classification are not possible for Spec2im, as the spectra required for a forward pass are not available for unmeasured points. In this case, stochasticity can be simply introduced by randomly sampling points with some probability.
Figure 5: Trial of Im2spec Curiosity Algorithm on pre-acquired PFM data: (a) Polarization ground-truth image, (b) Curiosity algorithm exploration path, and (c) a ground-truth hysteresis loop and corresponding Im2spec prediction. (d) Im2spec MSE error, (e) Predicted error, and (f) Exploration reward after final measurement iteration. Scale bar in the images indicate a length of 100 nm.
Figure 6: Minimum loss achieved by Im2spec with Curiosity Algorithm vs Random sampling.
Figure 7: MSE for ten highest error points achieved by Im2spec with MC Dropout Curiosity Algorithm vs Random sampling.
Figure 8: Trial of Curiosity Algorithm real-time on PFM microscope for Spec2im and Im2spec. (a) Exploration path of Spec2im Curiosity Algorithm, (b) predicted Spec2im error after final measurement iteration, (c) a polarization ground-truth image-patch, and (d) corresponding Spec2im prediction. (e) Exploration path of Im2spec Curiosity Algorithm, (f) predicted Im2spec error after final measurment iteration, (g) a ground-truth hysteresis loop and corresponding Im2spec prediction. Scale bar in the images indicate a length of 100 nm.
Another difficulty is the fact that as Im2spec/Spec2im trains, the MSE values change rapidly. As a result of this non-stationary problem, it is very challenging to train an accurate error predictor. Since the errors decrease on average, the problem can be made more stationary by training the error predictor on the errors divided by the mean error. These normalized MSE errors change much more slowly as Im2spec/Spec2im trains, and allow the error predictor to only account for relative changes in MSE error. It should be noted that even with this modification, the error predictor required a large learning rate and multiple epochs of training after each measurement in order to keep up with the changing errors.
We tested the Im2spec curiosity algorithm on the aforementioned pre-acquired PFM spectroscopic dataset in order to quantitatively determine its effectiveness. The PFM Polarization image ($P=A\sin(\theta)$), where $A$ is the piezoresponse amplitude and $\theta$ is the phase signal, is shown in Fig. 5(a). We benchmarked curiosity sampling based on predicted error and exploration reward against random sampling. To begin the algorithm, 30 initialization points were seeded, and the algorithm was then run for the next 170 points to sample based on the curiosity metric. The exploration path taken by the algorithm is shown in Fig. 5(b). It is evident that much of the sampling is occurring on the pre-exisiting domain walls, although several clusters of points within the domains are also sampled. The trained im2spec model after the 200 iterations appears to produce decent predictions compared with the ground truth, as shown in Fig. 5(c) for a chosen location. The MSE of im2spec is overall quite low, shown in Fig. 5(d) and does not appear spatially localized. The error predictor predicts maximal errors within the domains, and lowest errors at the domain walls, which also reflects the inverse of the sampled regions, as expected. The exploration reward, after the final measurement iteration, is mapped in Fig. 5(e) and again shows only a few isolated points with high errors. We benchmarked this against random sampling, and the results of the overall loss metrics after running 100 trials are shown in Fig. 6, and show clearly that the curiosity algorithm results in an overall lower loss than random sampling.
In addition, we tested a modified curiosity algorithm which, in addition to latent space exploration reward, samples based on Im2spec Monte-Carlo Dropout (MCD) uncertainty during the exploration phase. While the addition of MCD uncertainty did not directly improve Im2spec loss, it reduced the Im2spec MSE for the ten highest error points (Fig. 7). This behavior suggests that enhancing exploration with MCD helps train Im2spec on points with poorly understood structure property relationships, but are not abundantly represented in the sample data, as opposed to points with low error, but are highly represented in the sample data. It should be noted that one of the challenges of this algorithm is that there may exist points that continue to contain high errors regardless of the number of training data points, if there are minimal structure-property correlations in these points (for example, if there is only noise in these areas). For such instances, the algorithm should be modified to avoid trapping in these learning plateaus, ans strategies can include either direct human intervention, injected noise in the action space, or simple methods such as avoiding similar image patches to past samples if the loss is not decreasing beyond a simple threshold.
III.3 Real-time microscope deployment
Given these promising results on pre-acquired data, we moved to implement the curiosity algorithm on the microscope for real-time adaptive sampling. For additional difficulty, we changed the sample to one with a more complex domain structure, a thin film $PbTiO_{3}$ sample with a hierarchical domain strcuture that has been previously investigated liu2022exploring , and implemented the method using our AEcroscopy platform for microscope automation and acquisitionliu2024aecroscopy . Here, we tested both the spec2Im as well as the inverse, Im2spec, for the curiosity algorithm, and plot the results in Fig. 7. The exploration path the algorithm took for the Spec2Im case is shown in Fig. 8(a), and indicates a diverse spread of points across multiple different domain structures. Predicted errors are still spatially localized, but observing examples of predicted images compared with ground truth images show a decent predictive capability (Fig. 8(c,d)). The Im2Spec model shows a different exploration path, with many more points in the darker regions of the image, and the error map appears highly localized, potentially indicating that more points would need to be measured for more accurate modeling. Nevertheless, analysis of the actual predictions shows a decent corroboration with the ground truth (e.g., Fig. 8(g)).
IV Conclusions
In summary, we present two different workflows for curiosity driven spectral search. These frameworks utilize latent encodings for error prediction. While the first model utilizes latent space trained for spectral reconstruction, the second autoencoder model describes a generalized approach to train latent embeddings to predict spectral mismatch error. The curiosity algorithm was successful in sampling regions optimal for training Im2spec/Spec2im. On a preacquired dataset, we demonstrated that the curiosity algorithm outperformed random sampling. The algorithm was able to identify regions with complex structure-property relationships, particularly domain boundaries, and preferentially sample these regions in order to minimize Im2spec/Spec2im loss.
We implemented the workflow on a PFM microscope and found that the exploration paths optimizing Im2spec and Spec2im were different. This discrepancy is fundamentally caused by the in-existence of a bijection between domain structures and hysteresis loops. That is, several structures can produce the same hysteresis loop (for example, structures that are identical apart from a rotation). As a result, a single implementation of the curiosity algorithm is not sufficient for simultaneously optimizing both the forward and inverse problem. In practice, one must choose the algorithm better suited for the given application.
This curiosity based approach is a stepping stone to several novel autonomous microscopy workflows. For example, error prediction can be used to identify regions for which model error is high and does not decrease despite additional measurements, prompting more advanced spectroscopies to be performed in that region. Moreover, the convolutional neural networks may be replaced with theoretical models, in which case the curiosity algorithm would actively sample spectra for which the theory fails, offering insights informing new theoretical models.
V Data Availability Statement
The code and data used for model training and analysis can be found in the repository: https://github.com/cylindrical-penguin/Curiosity-Driven-RL-for-PFM.
VI Conflict of Interest
The authors have no conflict of interest to declare.
VII Acknowledgment
Algorithmic development was supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences, MLExchange Project, award number 107514. The experimental work was supported by the Center for Nanophase Materials Sciences (CNMS), which is a US Department of Energy, Office of Science User Facility at Oak Ridge National Laboratory. GN acknowledges support from QIS Infrastructure Project. HF acknowledges MEXT Program: Data Creation and Utilization Type Material Research and Development Project (Grant Number JPMXP1122683430). J.-C.Y. acknowledges the financial support from the National Science and Technology Council (NSTC), Taiwan, under grant numbers NSTC 112-2112-M-006-020-MY3 and 113-2124-M-006-010.